OpenStudy (anonymous):

A piece of cheese cut in the shape of a cube is 27 cubic inches. What is the length of each edge of the piece of cheese?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

x^3 = 27 x= sqrt(27) x=9

OpenStudy (anonymous):

um, no...9 * 9 * 9 is not 27

OpenStudy (anonymous):

its 3

OpenStudy (anonymous):

apologies!

OpenStudy (anonymous):

no - 3 is the answer - its cube root of 27

OpenStudy (anonymous):

suparfly, the question to ask your self, is, x * x * x = 27... what is x?

OpenStudy (anonymous):

It is 3.. \[\sqrt[3]{27} = 3\]... So the length of each side of the cube is 3 inches

OpenStudy (anonymous):

its 3 im right not negative three

OpenStudy (anonymous):

try a number, see if it is right, if not, re adjust your guess and try again.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so is it 3 inches or cubic inches

OpenStudy (anonymous):

3 inches, because the measurement of the edge is linear, for example you could measure it with a ruler. cubic is when you are talking about the actual volume of the object.

OpenStudy (anonymous):

ok